### v1

# Nonstrange and strange quark matter at finite temperature within modified NJL model and protoquark stars

Abstract
We extend the modified Nambu–Jona-Lasinio (NJL) model—incorporating exchange interactions via the Fierz transformation—to finite temperatures in both two- and three-flavor scenarios, and investigate the properties of protoquark stars in $\beta$-equilibrium. Our results show that increasing the strength of exchange interactions, characterized by the parameter $\alpha$, changes the chiral phase transition from first-order to crossover. We examine the effects of finite temperature, lepton fraction, and exchange interactions on the equation of state (EOS), finding that the EOS in the crossover regime is significantly stiffer than in the first-order case. In certain regions of parameter space within the crossover scenario, strange quark matter can support the existence of twin star configurations.
I Introduction
Neutron stars are the remnants of supernova explosions. They typically have masses in the range of $1-2\,M_{\odot}$, radii on the order of $\sim 10$ km, and are born with temperatures as high as $10^{11}$ K. Within a few days, they cool to approximately $10^{10}$ K by emitting neutrinos. In stellar modeling, the internal structure of a neutron star is determined by the assumed equation of state (EOS), which remains uncertain due to the incomplete understanding of the strong interaction under extreme conditions. At present, extensive studies have been conducted on proto-neutron stars with or without quarks at finite temperature Prakash et al. (1997); Reddy et al. (1998); Steiner et al. (2000); Pons et al. (2001); Steiner et al. (2001); Nicotra et al. (2006); Burgio & Plumari (2008); Yasutake & Kashiwa (2009); Shao (2011); Chen et al. (2012); Menezes et al. (2015); Raithel et al. (2019).
According to the Bodmer-Witten hypothesis Bodmer (1971); Witten (1984), compact stars can be composed entirely of deconfined quarks, forming self-bound quark stars. The composition of these self-bound stars hinges on a fundamental question—the true ground state of baryonic matter at zero pressure and temperature. After decades of theoretical speculation, both strange quark stars, composed of strange quark matter Chakrabarty et al. (1989); Dey et al. (1998); Chakrabarty (1993); Buballa & Oertel (1999); Peng et al. (1999); Wen et al. (2005); Li et al. (2010); Zhou et al. (2018); Li et al. (2018); Xia et al. (2019); Peng et al. (2000); Bai & Liu (2021); Xia et al. (2021); Yuan et al. (2022); Zhou et al. (2024); Zhang et al. (2024), and up-down quark stars, containing only up and down quark matter Holdom et al. (2018); Zhao et al. (2019); Ren & Zhang (2020); Zhang (2020); Zhang & Mann (2021); Restrepo et al. (2023); Su et al. (2024); Yang et al. (2025), have emerged as viable alternative models to neutron stars. Potential candidates for quark stars include compact objects with gravitational masses $\lesssim 1\,M_{\odot}$, such as PSR J1645–0317 (PSR B1642–03) Taylor et al. (1993), PSR J1830–1059 (PSR B1828–11) Taylor et al. (1993), 4U 1728–34 Li et al. (1999), and the X-ray bursters GRO J1744–28 Cheng et al. (1998) and SAX J1808.4–3658 Li et al. (1999). A recent analysis of the supernova remnant HESS J1731-347 suggests that its central compact object has a remarkably low mass of $M=0.77^{+0.20}_{-0.17}\,M_{\odot}$ and a radius of $R=10.4^{+0.86}_{-0.78}$  km Doroshenko et al. (2022), reigniting interest in the possible existence of self-bound quark stars. While observations of isolated compact stars provide constraints on the EOS of cold dense matter, many dynamical phenomena–such as core-collapse supernovae, the formation and cooling of proto-compact stars, and compact star mergers–depend on the EOS at finite temperatures. Thus, further investigation into the thermal effects on quark matter and the evolution of proto-quark stars is both crucial and intellectually stimulating Menezes et al. (2006); Dexheimer et al. (2013, 2014); Chu et al. (2017, 2019).
Due to the nonperturbative challenges at finite baryon densities, effective models that incorporate key features of QCD offer a practical alternative. For example, Ref. Lopes et al. (2021) employed the modified MIT bag model to study the quark matter at finite temperature and hot stars. To effectively capture the dynamics of spontaneous chiral symmetry breaking in QCD, the Nambu–Jona-Lasinio (NJL) model 1992RvMP…64..649K ; Hatsuda & Kunihiro (1994); 2005PhR…407..205B serves as a particularly suitable choice, which successfully reproduces the spectrum of low-lying mesons 2005PhR…407..205B ; 1992RvMP…64..649K ; Hatsuda & Kunihiro (1994), and has been widely applied and extended to describe quark matter in compact star physics (see, e.g., Refs. Huang et al. (2003); Blaschke et al. (2005, 2005); Wang et al. (2019); Minamikawa et al. (2021); Xia (2024); Gholami et al. (2025, 2024) as incomplete lists). Within this framework, Ref. Chu et al. (2017) investigates the role of isovector interactions and the behavior of the symmetry free energy of quark matter at finite temperature. In Ref. Menezes et al. (2006), the authors compare the predictions of the NJL model and the MIT bag model for quark matter at finite temperature and the properties of quark stars.
To improve the description of strongly interacting matter at high baryon chemical potentials, a modified version of the NJL model has been proposed in Ref. Wang et al. (2019). This extended NJL model incorporates both the original NJL Lagrangian and its Fierz-transformed counterpart, with their respective contributions weighted by the $1-\alpha$ and $\alpha$ parameters. This modified NJL model has been successfully applied in various studies, including the study of the critical endpoint in the QCD phase diagram Yang et al. (2019); Yu et al. (2020), color superconductivity Su et al. (2020); Yuan et al. (2023); Yuan & Li (2024); Yang et al. (2025), as well as the properties of cold quark matter and quark stars Wang et al. (2020); Yuan et al. (2022); Su et al. (2024). In this work, as an extension to previous studies, we first employ this modified NJL-type model to describe two-flavor and three-flavor quark matter at finite temperature, both with and without trapped neutrinos. We then systematically investigate the impact of exchange interactions on the properties of chiral phase transition and explore their implications for both strange and nonstrange protoquark stars.
This paper is organized as follows. In Sec. II, we will present the modified NJL model employed in this work to describe the nonstrange and strange quark matter at finite temperature, including the Fierz transformed interactions. Section III and Sec. V discuss the results on quark matter EOS at finite temperature and proto-quark stars. Our results are summarized in Sec. VI.
II The modified NJL models
In this section, we introduce the modified NJL model incorporating exchange channels via Fierz transformations, which account for the effects of rearranging fermion field operators.
II.0.1 two-flavor modified NJL model
The Lagrangian of the two-flavor NJL model reads:
$\mathcal{L}_{\mathrm{NJL}}^{~2f}=$ $\mathcal{L}_{0}+G\left[(\bar{\psi}\psi)^{2}+\left(\bar{\psi}i\gamma^{5}\tau\psi\right)^{2}\right]\ ,$ (1)
$\mathcal{L}_{0}=$ $\bar{\psi}\left(i\gamma^{\mu}\partial_{\mu}-m+\mu\gamma^{0}\right)\psi\ .$
Here, $\mathcal{L}_{0}$ denotes the relativistic free field, $\mathcal{L}_{0}=\bar{\psi}\left(i\gamma^{\mu}\partial_{\mu}-m+\mu\gamma^{0}\right)\psi$, which describes the propagation of non-interacting fermions. $G$ is the four-fermion interaction coupling constant. $\psi$ is the quark field operator with color, flavor, and Dirac indices. $\mu$ is the flavor-dependent quark chemical potential. The quark mass matrix $m=\mathrm{diag}(m_{u},m_{d})$ is diagonal in flavor space and includes the small current quark masses, thereby introducing a slight explicit breaking of chiral symmetry.
For further considering the effect of a rearrangement of fermion field operators, we apply the Fierz transformation to the interaction terms in the NJL models, as discussed in Refs. Wang et al. (2019); Yang et al. (2019); Yu et al. (2020); Su et al. (2020); Yuan et al. (2023); Yuan & Li (2024); Wang et al. (2020); Yuan et al. (2022); Su et al. (2024). As a purely technical device to examine the exchange channels influence that occur in quartic products at the same space-time point 1992RvMP…64..649K ; 2005PhR…407..205B , the Fierz identity of the four-fermion interactions in the two-flavor NJL model is
$\mathcal{F}(\mathcal{L}_{\sigma}^{4})=$ $\frac{G}{24}\left[2(\bar{\psi}\psi)^{2}+2\left(\bar{\psi}i\gamma^{5}\tau\psi\right)^{2}-2(\bar{\psi}\tau\psi)^{2}\right$. (2)
$-2\left(\bar{\psi}i\gamma^{5}\psi\right)^{2}-4\left(\bar{\psi}\gamma^{\mu}\psi\right)^{2}-4\left(\bar{\psi}i\gamma^{\mu}\gamma^{5}\psi\right)^{2}$
$\left.+\left(\bar{\psi}\sigma^{\mu\nu}\psi\right)^{2}-\left(\bar{\psi}\sigma^{\mu\nu}\tau\psi\right)^{2}\right]\ .$
We note that we only consider the contribution of color singlet terms for simplicity. One can see that, through the Fierz transformation, all exchange interaction channels of the original Lagrangian are released. In Eq. (2), the Fierz transformed Lagrangian contains not only the scalar and pseudoscalar interactions, but also vector and axialvector interaction channels.
Due to the mathematical equality between the original interactions and Fierz-transformed interactions, we can combine them using a weighting factor $\alpha$. The factor $\alpha$ reflects the competition between the original interaction channels and the exchange interaction channels. Then the effective Lagrangian becomes
$\mathcal{L}_{\rm eff}^{~2f}=\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m+\mu\gamma^{0})\psi+(1-\alpha)\mathcal{L}_{\mathrm{int}}^{~2f}+\alpha\mathcal{F}(\mathcal{L}_{\mathrm{int}}^{~2f})\ .$ (3)
When $\alpha=0$, the effective Lagrangian reduces to the standard NJL model. Unlike the conventional approach of artificially adding the vector interaction to the standard NJL Lagrangian when studying systems at finite chemical potential, the modified Lagrangian in Eq. (3) introduces vector interactions self-consistently Wang et al. (2019).
Under the mean-field approximation, the scalar and vector interactions in the Fierz-transformed channels in Eq. (2) also contribute at finite densities, thereby modifying the effective quark mass $M$ and the effective quark chemical $\tilde{\mu}$. The corresponding mass gap equation and the effective chemical potential for each quark are given as follows:
$M=$ $m-2\left[(1-\alpha)+\frac{\alpha}{12}\right]G\sum_{f=u,d}\sigma_{f}$ (4)
$=$ $m-2G^{\prime}\sum_{f=u,d}\sigma_{f}\ ,$
$\tilde{\mu}$= $\mu-\frac{\alpha}{3}G\sum_{f=u,d}\rho_{f}\ ,$
where we define $G^{\prime}=(12-11\alpha)G/12$. The quark condensate $\langle\bar{\psi}\psi\rangle$ and quark number density $\left\langle\psi^{+}\psi\right\rangle$ are denoted as $\sigma$ and $\rho$, respectively, which are the average values of operators, $\bar{\psi}\psi$ and $\psi^{+}\psi$, in the ground state.
In finite-temperature field theory Kapusta and Gale (2011), the linearization of $\mathcal{L}_{\rm eff}$ in the vicinity of the expectation values and the application of Matsubara formalism yield the thermodynamical potential per volume of quark matter as follows:
[…]
$-6\sum_{f=u,d}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\left[E_{f}+T\ln\left(1+e^{-\left(E_{f}+\tilde{\mu}_{f}\right)/T}\right)\right$.
[…]
$-\frac{\alpha G}{6}\left(\rho_{u}{}^{2}+\rho_{d}{}^{2}\right)+\Omega_{0}\ .$
From the grand canonical potential density in Eq. (5), all thermodynamic quantities of interest can be calculated using standard thermodynamic relations. The pressure is
$P=-\Omega$ (6)
$=6\sum_{f=u,d}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\left[E_{f}+T\ln\left(1+e^{-\left(E_{f}+\tilde{\mu}_{f}\right)/T}\right)\right$.
[…]
$+\frac{\alpha G}{6}\left(\rho_{u}{}^{2}+\rho_{d}{}^{2}\right)-\Omega_{0}\ ,$
and the energy density is
${\color[rgb]{0.0,0.0,0.8}\epsilon}=-P+TS+\sum_{f}\mu_{f}n_{f}$ (7)
$=6\sum_{f=u,d}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\left[E_{f}\left(\bar{n}_{f}+n_{f}-1\right)\right]$
$+2G^{\prime}\left(\sigma_{u}{}^{2}+\sigma_{d}{}^{2}\right)+\frac{\alpha G}{6}\left(\rho_{u}{}^{2}+\rho_{d}{}^{2}\right)+\Omega_{0}\ ,$
with
$n_{f}\left(T,\tilde{\mu}\right)=\frac{1}{1+\exp\left(\frac{E_{f}-\tilde{\mu}}{T}\right)}\ ,$ (8)
$\bar{n}_{f}\left(T,\tilde{\mu}\right)=\frac{1}{1+\exp\left(\frac{E_{f}+\tilde{\mu}}{T}\right)}\ ,$
which is the Fermi-Dirac distribution with $E_{f}=\sqrt{p^{2}+M_{f}^{2}}$.
II.0.2 (2+1)-flavor modified NJL model
The Lagrangian density of the (2+1)-flavor NJL model is given by
$\mathcal{L}_{\mathrm{NJL}}^{~3f}$ $=\mathcal{L}_{0}+\mathcal{L}_{\mathrm{int}}^{~3f}\ ,$ (9)
$\mathcal{L}_{\mathrm{int}}^{~3f}$ $=\mathcal{L}_{\sigma}^{4}+\mathcal{L}_{\sigma}^{6}\ ,$
where $\mathcal{L}_{\sigma}^{4}$ and $\mathcal{L}_{\sigma}^{6}$ represent the four-fermion and six-fermion interaction terms, respectively. These interaction terms are expressed as
$\mathcal{L}_{\sigma}^{4}$ $=\sum_{i=0}^{8}G\left[\left(\bar{\psi}\lambda_{i}\psi\right)^{2}+\left(\bar{\psi}i\gamma^{5}\lambda_{i}\psi\right)^{2}\right]\ ,$ (10)
$\mathcal{L}_{\sigma}^{6}$ $=-K\left(\operatorname{det}\left[\bar{\psi}\left(1+\gamma^{5}\right)\psi\right]+\operatorname{det}\left[\bar{\psi}\left(1-\gamma^{5}\right)\psi\right]\right)\ .$
Here, $G$ and $K$ denote the coupling constants for the four-fermion and six-fermion interactions, respectively. The matrices $\lambda_{i}\;(i=1\rightarrow 8)$ correspond to the Gell-Mann matrices in flavor space, while $\lambda_{0}=\sqrt{2/3}\;I_{0}$, with $I_{0}$ being the identity matrix. Applying the Fierz transformation to the four-fermion scalar and pseudoscalar interaction terms, considering only the contributions from color-singlet channels, yields
$\mathcal{F}(\mathcal{L}_{\sigma}^{4})=-\frac{G}{2}\left[\left(\bar{\psi}\gamma_{\mu}\lambda_{i}^{0}\psi\right)^{2}-\left(\bar{\psi}\gamma_{\mu}\gamma_{5}\lambda_{i}^{0}\psi\right)^{2}\right]\ .$ (11)
Since the Fierz transformation of the six-fermion interaction is defined to preserve invariance under all possible permutations of the quark spinors $\psi$ appearing in the interaction 1992RvMP…64..649K , the six-fermion term remains unchanged:
$\mathcal{F}(\mathcal{L}_{\sigma}^{6})=\mathcal{L}_{\sigma}^{6}\ .$ (12)
Thus, the effective Lagrangian takes the form
$\mathcal{L}_{\rm eff}^{~3f}=\bar{\psi}(i\gamma^{\mu}\partial_{\mu}-m+\mu\gamma^{0})\psi+(1-\alpha)\mathcal{L}_{\mathrm{int}}^{~3f}+\alpha\mathcal{F}(\mathcal{L}_{\mathrm{int}}^{~3f})\ .$ (13)
In the mean-field approximation, the mass gap equations and the effective chemical potential $\mu_{f}^{*}$ for a given flavor $f=i,j,k$ are given by
$M_{i}$ $=m_{i}-4(1-\alpha)G\sigma_{i}+2K\sigma_{j}\sigma_{k}$ (14)
$=m_{i}-4G^{\prime}\sigma_{i}+2K\sigma_{j}\sigma_{k},$
$\tilde{\mu}_{f}$ $=\mu_{f}-\frac{2}{3}\alpha G\sum_{f^{\prime}=u,d,s}\rho_{f^{\prime}}\ ,$
where we define $G^{\prime}=(1-\alpha)G$, and $i,j,k$ represent even permutations of $u,d,s$. From Eq. (14), it is evident that the introduction of the Fierz-transformed interactions contributes to the effective chemical potential and modifies the gap equations.
The grand canonical potential density of the interacting strange quark matter system is
[…]
$-6\sum_{f=u,d,s}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\left[E_{f}+T\ln\left(1+e^{-\left(E_{f}+\tilde{\mu}_{f}\right)/T}\right)\right$.
[…]
$-\frac{\alpha G}{3}\left(\rho_{u}+\rho_{d}+\rho_{s}\right)^{2}-4K\sigma_{u}\sigma_{d}\sigma_{s}+\Omega_{0}\ .$
Then all thermodynamic quantities of interest can be calculated. The pressure is
$P=-\Omega$ (16)
$=6\sum_{f=u,d,s}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\left[E_{f}+T\ln\left(1+e^{-\left(E_{f}+\tilde{\mu}_{f}\right)/T}\right)\right$.
[…]
$+\frac{\alpha G}{3}\left(\rho_{u}+\rho_{d}+\rho_{s}\right)^{2}+4K\sigma_{u}\sigma_{d}\sigma_{s}-\Omega_{0}\ ,$
and the energy density is
$\epsilon=-P+TS+\sum_{f}\mu_{f}n_{f}$ (17)
$=6\sum_{f=u,d,s}\int\frac{\mathrm{d}^{3}p}{(2\pi)^{3}}\left[E_{f}\left(\bar{n}_{f}+n_{f}-1\right)\right]$
$+2G^{\prime}\left(\sigma_{u}{}^{2}+\sigma_{d}{}^{2}+\sigma_{s}{}^{2}\right)+\frac{\alpha G}{3}\left(\rho_{u}+\rho_{d}+\rho_{s}\right)^{2}$
$-4K\sigma_{u}\sigma_{d}\sigma_{s}+\Omega_{0}\ .$
Here, we choose the constant in Eq. (15) such that $P$ and $\epsilon$ vanish in vacuum, i.e., we choose $\Omega_{M}(0,0)=0$.
So far, we have successfully derived the thermodynamic potential densities at finite chemical potential and temperature. In order to ensure thermodynamic consistency, the condensates must be obtained through appropriate differentiation of the thermodynamic potential. The self-consistent solutions are those that correspond to the stationary points of the potential, which are defined by:
$\frac{\delta\Omega}{\delta\sigma}=\frac{\delta\Omega}{\delta\tilde{\mu}}=0\ ,$ (18)
from which we obtain the quark condensate expressed as
$\sigma_{f}=-\frac{N_{c}M_{f}}{\pi^{2}}\int_{0}^{\Lambda}\frac{p^{2}}{E_{f}}\mathrm{d}p\left[1-n_{f}(T,\tilde{\mu}_{f})-\bar{n}_{f}(T,\tilde{\mu}_{f})\right]\ ,$ (19)
and the quark number densities are
[…]
for $f=u,d,(s)$.
III Quark matter in $\beta$-equilibrium
A newborn neutron star forms in the aftermath of a successful supernova explosion, when the stellar remnant becomes gravitationally decoupled from the expanding ejecta and reaches core temperatures as high as $\sim 50\;\text{MeV}$ Lattimer & Prakash (2004); Weber (2005). In this work, we consider two distinct scenarios: an early stage in which neutrinos are trapped within the stellar interior, and a later stage after deleptonization, where neutrinos have escaped and their chemical potential vanishes. Although an isentropic description is generally considered more realistic for modeling proto-quark stars, we adopt an isothermal approach, as made in previous work Santos & Menezes (2004); Yasutake & Kashiwa (2009); Lu et al. (2019); Lopes et al. (2021); Pal & Chaudhuri (2024), to enable direct comparisons across different temperature cases.
In $\beta$-equilibrated matter, the contribution of leptons is included as a free Fermi gas in the calculation of the energy density, pressure, and entropy density. The relations among the chemical potentials of the various particle species without neutrinos under $\beta$-equilibrium are given by
$\mu_{s}=\mu_{d}=\mu_{u}+\mu_{e}\ .$ (21)
One has to require also the charge-neutrality of the quark matter,
$\frac{2}{3}\rho_{u}-\frac{1}{3}\rho_{d}-\frac{1}{3}\rho_{s}-\rho_{e}=0\>,$ (22)
and the baryon number conservation,
$\frac{1}{3}\left(\rho_{u}+\rho_{d}+\rho_{s}\right)=\rho_{\rm B}\>,$ (23)
is satisfied with $\rho_{\rm B}$ being the baryon number density. The baryon chemical potential satisfies $\mu_{\rm B}=\mu_{u}+\mu_{d}+\mu_{s}$ for strange quark matter, and $\mu_{\rm B}=\mu_{u}+2\mu_{d}$ for nonstrange quark matter.
In earlier stages, when the neutrinos are still trapped in the interior of the star, Eq. (21) is replaced by
$\mu_{s}=\mu_{d}=\mu_{u}+\mu_{e}-\mu_{\nu e}\ .$ (24)
and the lepton contribution $Y_{l}$ is set to be constant
$(\rho_{e}+\rho_{\nu e})/\rho_{B}=Y_{l}\ .$ (25)
The lepton has no interactions, so the grand canonical potential density is the free grand canonical potential density. The lepton densities are given by
$\rho(\mu_{e},T)=$ $2\int\frac{p^{2}\mathrm{d}p}{2\pi^{2}}\left[n_{e}(T,\mu)-\bar{n}_{e}(T,\mu)\right]\ ,$ (26)
$\rho(\mu_{\nu e},T)=$ $\int\frac{p^{2}\mathrm{d}p}{2\pi^{2}}\left[n_{e}(T,\mu)-\bar{n}_{e}(T,\mu)\right]\ .$
The EOS of dense warm matter is obtained by including both the contribution of quarks and leptons, namely the total energy density, pressure and entropy density are given by
[…]
Figure 1: Upper panel: The effective quark masses of u,d,su,\ d,\ s quarks versus quark chemical potential $\mu$ both for strange quark matter and nonstrange quark matter in the absence of neutrinos ($Y_{l}=0.0$), under various values of $\alpha$ and temperature $T$. The results for several representative parameter sets are shown: ($\alpha=0.0,\ T=0.04$ GeV), ($\alpha=0.0,\ T=0.00$ GeV), ($\alpha=0.5,\ T=0.00$ GeV), ($\alpha=0.8,\ T=0.00$ GeV), ($\alpha=0.9,\ T=0.00$ GeV). Middle panel: The corresponding results for quark number densities of u,d,su,\ d,\ s quarks as functions of $\mu$. Lower panel: The electron chemical potential $\mu_{e}$ as a function of $\mu$.
Figure 2: The effective quark masses $M_{u}$, $M_{d}$, and $M_{s}$, the electron chemical potential $\mu_{e}$, and the $d$ quark chemical potential $\mu_{d}$ are shown as functions of the $u$ quark chemical potential in the neutrino-trapped scenario, for a representative case with $\alpha=0.5$, $Y_{l}=0.1$, and $T=0.04$ GeV. Results are presented for both strange and nonstrange quark matter.
IV quark matter properties at finite temperature
IV.1 fitting parameters
Before doing the calculation, we should fit the NJL model parameters. Since the NJL model is not renormalizable, a momentum cutoff $\Lambda$ should be introduced to regulate the ultraviolet divergences. Here, we use the three momentum cutoff. After defining the new coupling constant $G^{\prime}$, and keeping the expression of gap equation the same as the widely used one in Eqs. (4) and (14), at zero temperature and chemical potential, apart from $\alpha$, the fixing of the model parameters is the same with the original version of the NJL model Hatsuda & Kunihiro (1994). Similar to the procedure in Ref. Hatsuda & Kunihiro (1994), after fixing the masses of the up and down quarks by equal values, the other parameters $m_{s},\Lambda_{\mathrm{UV}},G^{\prime}$, $K$ are chosen to reproduce the experimental data of the pion decay constant and pion mass. The employed model parameters are shown in Table 1.
Table 1: NJL model parameters are shown. The units of the coupling constants $G^{\prime}$ and $K$ are $\mathrm{MeV}^{-2}$ and $\mathrm{MeV}^{-5}$, respectively, and the other parameters have the units of $\mathrm{MeV}$.
$m_{u}$ $m_{s}$ $\Lambda_{\mathrm{UV}}$ $G^{\prime}$ $K$
Two-flavor $5.0$ / $653.0$ $4.930\times 10^{-6}$ /
$(2+1)$-flavor $5.5$ 104.7 602.3 $5.058\times 10^{-6}$ $1.559\times 10^{-13}$
To obtain the EOS, the following set of equations must be solved self-consistently. For neutrino-free matter, the relevant system consists of the gap equations of the NJL model in Eqs. (4) and (14), along with the conditions for $\beta$-equilibrium in Eq. (21) and charge neutrality in Eq. (22). These constraints reduce the number of independent chemical potentials to one. We choose the up quark chemical potential $\mu_{u}$ as the input quantity, with the remaining chemical potentials, $\mu_{d}$, $\mu_{s}$ and $\mu_{e}$ determined as functions of $\mu_{u}$. In the case of neutrino-trapped matter, the system includes the same NJL gap equations, together with the modified $\beta$ equilibrium condition in Eq. (24), charge neutrality in Eq. (22), and a fixed lepton fraction in Eq. (25). Again, only one independent chemical potential remains.
IV.2 without neutrinos
In this section, we investigate the effects of exchange interactions and temperature on the properties of quark matter under conditions of $\beta$-equilibrium and charge neutrality, neglecting neutrino contributions. Figure 1 presents the effective quark mas and quark number densities of u,d,su,\ d,\ s quarks, as well as electron chemical potential calculated within the $(2+1)$-flavor and two-flavor NJL models. At zero temperature, when $\tilde{\mu}_{f}<M_{f}$, the quark condensate is independent of the quark chemical potential, and the dynamical quark mass remains at its vacuum value. When $\tilde{\mu}_{f}>M_{f}$, the chiral symmetry begins to restore. As the parameter $\alpha$ becomes nonzero, vector interactions are introduced self-consistently into the system. Due to the repulsive nature of these interactions in the exchange channel, the dynamical quark mass decreases more gradually with increasing chemical potential $\mu$, particularly for the $s$ quark. This effect is clearly illustrated by comparing the cases of $\alpha=0.0,\ T=0.00$ GeV and $\alpha=0.5,\ T=0.00$ GeV. A further increase of $\alpha$ to 0.8 (0.9) enhances the role of vector interactions in the exchange channels, further suppresses chiral symmetry restoration, and drives the first-order chiral phase transition toward a crossover. In comparison with the NJL model without vector interactions ($\alpha=0.0,\ T=0.00$ GeV), increasing the temperature makes the quark condensate begins to restore at a relatively lower chemical potential.
As expected, at zero temperature, the quark number density stays zero when $\mu$ is smaller than the constituent quark mass $M_{u,d}$ and $M_{s}$. Once $\mu$ exceeds the corresponding dynamical quark mass, the quark number density exhibits a discontinuous jump to a nonzero value, signaling a first-order phase transition. In the two-flavor quark matter system, the restoration of the quark condensate leads to an increase in the up and down quark number densities, $\rho_{u,d}$. Due to the charge neutrality condition, this necessitates an increase in the electron fraction, and thus, the electron chemical potential $\mu_{e}$ rises with increasing $\mu$. In contrast, the behavior of $\mu_{e}$ in strange quark matter is more intricate. Increase the value of $\alpha$ enhances the vector interaction strength, and these repulsive interactions suppress the restoration of the strange quark condensate, thereby reducing the strange quark number density $\rho_{s}$. Consequently, particularly for $\mu\gtrsim 400$ MeV, a larger value of $\alpha$ leads to a higher electron chemical potential, compensating for the reduced negative charge from strange quarks in order to maintain overall charge neutrality. In particular, for $\alpha=0.9$, the strange quark condensate is rarely restored, yielding a system with a negligible strange quark population and, consequently, a significantly elevated $\mu_{e}$ at high $\mu$.
IV.3 neutrino trapped
To investigate the properties of proto-quark stars, we assume that neutrinos are trapped in the matter. In the following analysis, we consider a temperature range of $T=0-40\,\mathrm{MeV}$ and a total lepton fraction of $Y_{l}=0.1-0.4$, consistent with conditions expected in compact stars during the core-collapse phase of massive stellar evolution Burrows & Lattimer (1986). We will demonstrate that the strength of exchange interactions and the lepton fraction plays a more significant role than temperature in determining the stiffness of the EOS. The underlying reasons for this behavior will be discussed later.
For neutrino-trapped matter, we solve Eqs. (4 or 14), (22), (24), and (25) self-consistently, taking the chemical potential of the $u$ quark as the input quantity. The resulting physical solutions for two-flavor and (2+1)-flavor quark matter, corresponding to a representative case with $\alpha=0.5$, $Y_{l}=0.1$, and $T=0.04~\mathrm{GeV}$, are shown in Fig. 2. For (2+1)-flavor quark matter, When $\mu>220~\mathrm{MeV}$, the $u$ quark condensate begins to melt slightly, accompanied by the appearance of a small population of $u$ and $d$ quarks. Owing to flavor-mixing effects and the associated reduction of $M_{u}$ and $M_{d}$, the dynamical mass of the $s$ quark also decreases slightly. At $\mu=351~\mathrm{MeV}$, a first-order phase transition occurs, marked by a discontinuous drop in the dynamical masses $M_{u}$, $M_{d}$, and $M_{s}$. The effective $u$ quark mass differs slightly from the $d$ quark mass due to the $\beta$-equilibrium. For $\mu>365~\mathrm{MeV}$, the $ud$ quark masses are fully restored to their current masses, and the flavor-mixing effect becomes negligible. As the quark chemical potential increases further, the strange quark condensate continues to decrease, and the number density of $s$ quarks increases. In contrast, the two-flavor case exhibits a relatively low dynamical quark mass of approximately $311~\mathrm{MeV}$, generated via spontaneous chiral symmetry breaking, with only a weak discontinuity in the $M(\mu)$ relation. Due to the absence of strange quarks, both the electron chemical potential $\mu_{e}$ and the neutrino chemical potential $\mu_{\nu_{e}}$ increase monotonically with $\mu_{u}$, driven by the requirements of charge neutrality and $\beta$-equilibrium.
In Fig. 3, we examine the effects of lepton fraction $Y_{l}$, temperature $T$, and the strength of exchange interactions—characterized by the parameter $\alpha$—on both strange and nonstrange quark matter. We present the effective quark masses $M_{u}$ and $M_{s}$ as functions of the quark chemical potential $\mu$, along with the pressure as a function of baryon number density $\rho_{B}$ for various parameter sets. In the lower panel of Fig. 3, for the first-order chiral phase transition, we consider only the chirally restored quark matter phase predicted by the NJL model to be reliable; therefore, the results are shown starting from $2\rho_{0}$. Under the same conditions of $Y_{l}=0.1$ and $T=0.04$ GeV, comparing stronger vector interactions with $\alpha=0.5$ to those with $\alpha=0.0$ shows that the former shifts the onset of the first-order phase transition to higher quark chemical potentials and significantly suppresses the restoration of effective quark masses at high chemical potentials, especially that of the strange quark. A similar trend is observed in the case of nonstrange quark matter. As shown in the lower panel, the EOS becomes stiffer for larger $\alpha$, reflecting a stronger contribution from vector interactions in the exchange channels, for both strange and nonstrange quark matter. Compared to the case of $\alpha=0.8$, $Y_{l}=0.1$, and $T=0.01$ GeV, increasing the temperature to $T=0.04$ GeV accelerates the melting of the dynamical quark masses before $\mu\sim 400\;\text{MeV}$, causing the strange quark phase to emerge at lower chemical potentials. The similar behavior can also be found for nonstrange quark matter. The temperature has only a slight effect on the stiffness of the EOS, as can be seen from the $P(\rho_{B})$ relations. For fixed $\alpha=0.8$ and $T=0.04$ GeV, increasing the lepton fraction from $Y_{l}=0.1$ to $Y_{l}=0.4$ suppresses the restoration of dynamical masses and stiffens the EOS of quark matter. A higher lepton fraction implies a larger electron population, which, due to the charge neutrality condition, suppresses the abundance of other negatively charged particles, such as the $s$ quark. As a result, a higher $Y_{l}$ leads to a reduced fraction of $s$ quarks, thereby hindering the restoration of their dynamical mass.
Figure 3: The effective quark masses of u,d,su,\ d,\ s quarks versus quark chemical potential $\mu$ for various choices of the (2+1)-flavor and two-flavor NJL model parameters with neutrino trapped scenario for the study of their effects (see text for details). Lower panel: The pressure as a function of baryon number density, expressed in units of the nuclear saturation density $\rho_{0}$, for both strange quark matter and nonstrange quark matter.
V Proto-quark star properties
[…]
Figure 5: Left panel: The pressure versus the energy density (EOS) for several parameter sets calculated from the two-flavor modified NJL model. The EOSs are all smooth here. Right panel: Corresponding mass-radius relations for the self-bound nonstrange quark matter EOS for the six parameter sets shown in the left panel. The available astronomical constraints shown here are the same as in Fig. 4.
In the following, we examine the influence of the exchange interaction parameter $\alpha$, lepton fraction $Y_{l}$, and temperature $T$ on the EOS and the structure of protoquark stars. As shown in Sec. IV, increasing the strength of the exchange channel $\alpha$ alters the nature of the chiral phase transition from first-order to a crossover. Since the relationship between chiral symmetry restoration and deconfinement remains unclear, in this work, we assume that deconfinement occurs simultaneously with the chiral phase transition. Therefore, to effectively mimic confinement at low densities, a finite vacuum pressure is introduced. According to the properties of the chiral phase transition, we adopt different prescriptions for defining the vacuum pressure B-B:
  • For a first-order phase transition at low temperature and weak vector interaction, the chiral symmetry breaking phase (hadronic phase), which cannot be reliably described by the NJL model, will be transformed to the chiral restored phase (quark phase). Therefore, we consider only the quark phase after the transition and choose the minimum vacuum pressure $B_{0}$ such that the pressure vanishes at a finite energy density, as required for modeling self-bound quark stars. Any other choice of $B$ should be larger than $B_{0}$. For example, in Fig. 4, for the case of $\alpha=0.5$, $Y_{l}=0.4$, and $T=40~\mathrm{MeV}$, the chiral phase transition is of first order, resulting in a discontinuity in the energy density of the EOS. To ensure that the quark star is composed of chirally restored quark matter, the vacuum pressure must exceed $133,\mathrm{MeV}$, allowing the quark matter EOS to reach zero pressure at a finite energy density—thus ensuring the star is self-bound at the surface.
  • For a crossover chiral transition at relatively large temperature and strong vector interactions in the exchange channel, the transition is smooth. As a result, the choice of vacuum pressure is more flexible, and even a small vacuum pressure can ensure that the EOS yields a finite energy density at zero pressure.
We note that, according to the Bodmer–Witten hypothesis Bodmer (1971); Witten (1984), absolutely stable quark matter must have an energy per baryon lower than that of the $^{56}\mathrm{Fe}$ nucleus at zero pressure and zero temperature. Accordingly, for self-bound strange quark matter, the condition $(E/A){\mathrm{uds}}\leq 930~\mathrm{MeV}$ must be satisfied for the hypothesis of strange matter to hold. In contrast, bulk nonstrange quark matter should have an energy per baryon higher than that of the confined phase, i.e., $(E/A){\mathrm{ud}}\geq 930~\mathrm{MeV}$. For absolutely stable nonstrange quark matter, the stability condition is reversed: $(E/A){\mathrm{ud}}\leq 930~\mathrm{MeV}\leq(E/A){\mathrm{uds}}$. In our previous work using the modified NJL model with exchange interactions Yuan et al. (2022), we demonstrated that there is sufficient parameter space to support both absolutely stable nonstrange and strange quark matter at zero temperature. However, at finite temperature, these conditions are not necessarily required to be satisfied.
V.1 Strange proto-quark star properties
In Fig. 4, we present the EOS at finite temperature with neutrino trapping, along with the corresponding mass-radius relations for protoquark stars. The full EOSs represent results derived from the NJL model, including the chirally broken phase. To model self-bound quark matter that can form quark stars, a vacuum pressure $B$ is added to ensure the system reaches zero pressure at a finite density. For the case of $\alpha=0.5$, $Y_{l}=0.1$, and $T=10\,\mathrm{MeV}$, a discontinuous jump in energy density appears, characteristic of a first-order phase transition. In this scenario, the vacuum pressure must be at least $B_{0}=133\,\mathrm{MeV}$ to ensure that the pressure vanishes at the surface of the compact star. However, when the chiral phase transition becomes a crossover, the required vacuum pressure $B$ can be smaller than $B_{0}$.
In general, increasing the value of $\alpha$ enhances the repulsive vector interactions, leading to a stiffer EOS, which can be seen clearly in the comparison between $\alpha=0.5$ and $\alpha=0.8$. Previous studies have also shown that strong vector interactions are necessary to support massive compact stars 2005PhR…407..205B . For large values of $\alpha$, such as $\alpha=0.8$, the chiral phase transition becomes a crossover. This smooth transition typically results in a stiffer EOS compared to the case of a first-order transition. For the case of $\alpha=0.8,Y_{l}=0.1,\ B=133\;\text{MeV}$, increasing the temperature from $T=10\;\text{MeV}$ to $40\;\text{MeV}$ slightly softens the EOS, leading to a modest reduction in the maximum mass $M_{\mathrm{TOV}}$, while increasing the stellar radius at the same mass. Because the increase in energy exceeds the thermal contribution to pressure. A similar trend is also reported in Ref. Pal & Chaudhuri (2024), where, within the quasiparticle model incorporating vector interactions, the maximum mass decreases with increasing temperature in the isothermal scenario. When the total lepton fraction $Y_{l}$ increases from 0.1 to 0.4 (with $\alpha=0.8$, $T=0.01~\mathrm{MeV}$, and $B=100~\mathrm{MeV}$), the EOS exhibits a more complex behavior due to the intersection of the two curves. At higher densities, the EOS with a larger lepton fraction becomes stiffer, primarily due to the suppression of the strange quark population.
At $\alpha=0.8$, $Y_{l}=0.1$, and $T=10\;\text{MeV}$, a representative vacuum pressure of $B=110,\mathrm{MeV}$ yields a quark star model consistent with observational constraints from both the NICER mission 2019ApJ…887L..24M ; 2019ApJ…887L..21R ; 2021ApJ…918L..28M ; 2021ApJ…918L..27R and the compact object located in the supernova remnant HESS J1731–347 Doroshenko et al. (2022). Further reducing the vacuum pressure to $B=100\;\text{MeV}$ results in an even stiffer EOS, leading to a particularly intriguing outcome—the appearance of a twin-star configuration Alford et al. (2013); Alvarez-Castillo & Blaschke (2017); Christian et al. (2018); Montaña et al. (2019), i.e., two compact stars with nearly the same gravitational mass but substantially different radii, located on two separate stable branches of the mass–radius relation. This phenomenon can be attributed to the emergence of strange quarks at typical baryon number densities. As shown in Fig. 4, the appearance of strange quarks initially softens the EOS slightly, followed by a stiffening at higher densities. As discussed in Refs. Alford et al. (2013); Alvarez-Castillo & Blaschke (2017); Christian et al. (2018); Montaña et al. (2019), a sharp first-order phase transition inside compact stars can lead to twin-star configurations. In our study, we find that such configurations can also arise in some parameter sets even when the chiral transition is a smooth crossover. It should be noted that the gravitational wave constraints from GW170817 were derived from the late inspiral phase of the merger, under the assumption of zero temperature 2017PhRvL.119p1101A ; 2018PhRvL.121p1101A . Therefore, while we include these constraints for reference, they are not directly applicable to hot compact stars. The precise temperature of neutron stars during the inspiral phase remains an open question Laskos-Patkos et al. (2022). In the near future, we will investigate the influence of finite-temperature effects on tidal deformability and the I-Love-Q relations within the modified NJL models.
V.2 Nonstrange proto-quark star properties
Figure 5 shows the EOS of nonstrange quark matter derived from our modified NJL model, together with the corresponding mass–radius relations for several representative parameter sets. For both $\alpha=0.8$ and $\alpha=0.9$, the chiral phase transition is smooth, yielding continuous EOSs. To ensure that finite-temperature nonstrange quark matter is self-bound—i.e. capable of forming a quark star—we add a vacuum pressure term $B$. At $\alpha=0.8,\ T=10\;\text{MeV},\ B=80\;\text{MeV}$, raising the lepton fraction from $Y_{l}=0.1$ to $Y_{l}=0.4$ stiffens the EOS significantly, so that the resulting maximum masses are compatible with NICER constraints. Increasing the exchange channel strength from $\alpha=0.8$ to $\alpha=0.9$ likewise produces a marked stiffening of the EOS. Reducing the vacuum pressure from $B=80\;\text{MeV}$ to $B=60\;\text{MeV}$ not only stiffen the EOS but also affects the low-density regime, leading to larger stellar radii than in the $B=80\;\text{MeV}$ case.
Different with the case of (2+1)-flavor quark matter, increasing the temperature of nonstrange quark matter tends to stiffen the EOS, as the rise in pressure outweighs the thermal contribution to the energy density. For example, with $\alpha=0.8,Y_{l}=0.4,\ B=80\;\text{MeV}$, increase the temperature from $10\;\text{MeV}$ to $40\;\text{MeV}$ raises the maximum mass $M_{\rm TOV}$ from $1.93\,M_{\odot}$ to $2.05\,M_{\odot}$. The influence of temperature on nonstrange quark matter is analogous to that observed in neutron matter, leading to a stiffer EOS and an increased maximum mass at higher temperatures Koliogiannis & Moustakidis (2021); Wu et al. (2025). To remain consistent with current astronomical observations, a sufficiently strong vector interaction and a relatively small vacuum pressure are required. No parameter space is found in which proto-nonstrange quark stars can form a twin star configuration.
VI Summary
In this work, we investigate protoquark stars by combining the EOS of quark matter at finite temperature with a modified NJL model that incorporates exchange interactions, weighted by a parameter $\alpha$, for both two- and three-flavor cases. Unlike previous studies that introduce vector interactions manually, our approach uses the Fierz transformation to include scalar and vector contributions from the exchange channels self-consistently. Within this framework, we find that increasing the exchange interaction strength $\alpha$ modifies the chiral phase transition from a first-order transition to a crossover. Treating the vacuum pressure as a phenomenological parameter, we find that a first-order chiral phase transition within the NJL model necessitates a relatively large vacuum pressure to ensure vanishing pressure at the stellar surface, thereby yielding a softer EOS. An increase in the lepton fraction $Y_{l}$ stiffens the EOS for both strange and nonstrange quark matter at finite temperature. Temperature effects differ between strange and nonstrange quark matter: high temperatures tend to soften the EOS for strange quark matter, while relatively stiffening it for nonstrange quark matter. Notably, for protoquark stars composed of strange quark matter, we identify a region in parameter space that allows for the existence of twin star configurations.
Looking ahead, a more realistic treatment of protoquark stars should incorporate isotropic conditions and magnetic field effects Chu et al. (2020, 2022). Beyond the normal quark phase, the potential role of a color superconducting phase and pion condensation phase at finite temperature is also particularly intriguing—especially in scenarios where high baryon densities are realized in the early Universe due to a large lepton number Vovchenko et al. (2021); Ferreira et al. (2025). Within the same finite-temperature NJL framework, it would be highly interesting to investigate the influence of diquark condensates on the evolution of protoquark stars. The corresponding results would be crucial for understanding the cooling mechanism of the compact stars, and could shed light on the behavior of dense matter and its broader cosmological implications.
Acknowledgements.
We thank Chen Zhang, Pengcheng Chu, and Bingjun Zuo for their valuable comments and useful discussions. We also thank the PKU pulsar group for the helpful discussions. This work is supported by the National SKA Program of China (2020SKA0120100), the National Natural Science Foundation of China (Nos. 12003047, 12133003), the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0550300). Wen-Li Yuan is supported by the Special Funds of the National Natural Science Foundation of China (Grant No. 12447171) and the China Postdoctoral Science Foundation (Grant. No. 2025M773418). Professor Nobutoshi Yasutake is supported by the JSPS KAKENHI (Grant No. 24K07054).